2 cars A and B . A travelled at 7.00 with speed of 4 mile/h . B went after 30 minutes with speed of 8.5 mile/h. at what time will B reached to A .
Thanks alot
1
Expert's answer
2013-02-04T08:51:00-0500
When B reaches A, they travelled equal distance.
Distance that travelled A, is 4*t miles, where t is a timefor A to travel, in hours.
So, distance that was covered by B, is 8.5(t-0.5)
4*t = 8.5*(t-0.5)
solving that, we obtain that t is 17/18 hours = 56 + 2/3minutes
so, time when B will reach A is 7.57, when rounding.